# Global Compression Commander: Plug-and-Play Inference Acceleration for High-Resolution Large Vision-Language Models

### v4

@@ Abstract
Large vision-language models (LVLMs) excel at visual understanding and reasoning, but face efficiency challenges due to quadratic complexity in processing long multimodal contexts. While token compression techniques can reduce computational costs, existing approaches are designed for single-view LVLMs and fail to consider the unique multi-view characteristics of recent high-resolution LVLMs with dynamic tiling. While existing methods treat all tokens uniformly, our analysis reveals that global thumbnails can naturally guide the compression of local crops by providing holistic context for informativeness evaluation. In this paper, we first analyze dynamic tiling strategy comprehensively, revealing both the complementary nature between thumbnails and crops, and the distinctive characteristics across different crops. Based on our observations, we propose “Global Compression Commander” (i.e., GlobalCom2), a novel plug-and-play token compression framework for HR-LVLMs. GlobalCom2 leverages thumbnail as the “commander” to guide the compression process of local crops, adaptively preserving informative details while eliminating redundancy. Extensive experiments show that GlobalCom2 maintains over 90% performance while compressing 90% visual tokens, reducing FLOPs and peak memory to 9.1% and 60% respectively across multiple benchmarks. Our code is available at https://github.com/xuyang-liu16/GlobalCom2.
Figure 1: Design philosophy of “global-to-local” guided token compression. GlobalCom2 evaluates the information richness of crops from a global perspective to preserve informative regions while removing redundancy.
@@ 1 Introduction
By bridging visual encoders with Large Language Models (LLMs) [40, 35, 1], Large Vision-Language Models (LVLMs) [27, 28, 2, 24, 51] have achieved remarkable progress in multiple tasks. However, the large number of visual tokens leads to prohibitively expensive computation during inference [8, 38, 18, 13], especially for high-resolution images, limiting LVLMs’ practical applications.
To tackle the challenge of inference speed in LVLMs, researchers have primarily focused on model-centric approaches like knowledge distillation [4] and quantization [43], which typically require model retraining, consuming substantial computational resources thus limiting their practical applicability. Recent efforts have shifted towards data-centric approaches, exemplified by token compression methods [5, 47, 38, 31, 16], which consider the redundancy in LVLMs’ input visual signals to enhance inference speed by reducing processed tokens while preserving essential information. Most of these methods are architecture-agnostic and training-free, achieving optimal efficiency-accuracy trade-off while being complementary to model-centric approaches [8, 53, 52].
However, current token compression methods were designed for LVLMs processing single-view images (entire image without crop decomposition, as shown in Figure 1 top). With growing demand for high-resolution image understanding, advanced HR-LVLMs like LLaVA-NeXT [29] and InternVL 1.5 [9] adopt dynamic tiling, which divides an image into several crops encoded individually and concatenated with tokens from a global thumbnail (in Figure 1 top). While this multi-view approach enables more comprehensive visual understanding, it further increases the number of visual tokens and computational overhead.
Directly applying existing token compression methods to all views might seem straightforward, but such uniform compression overlooks two critical structural properties revealed by our analysis: (1) The inherent role of global thumbnails in extracting holistic context and guiding general vision understanding, demonstrated in Figure 2, and (2) The information richness disparity across crops quantified in Figure 3 (e.g., 2.4% average performance gap between discarding most/least informative crops). These oversights lead to over-compression of information-rich regions and disrupted visual-semantic hierarchy (see Figure 6).
To establish theoretically grounded compression principles for dynamic tiling-based HR-LVLMs, we first conduct systematic analysis to investigate the roles of global thumbnails and local crops in HR-LVLM image understanding capabilities. Through extensive empirical studies, we identify two key observations for token compression: Visual tokens from thumbnails and crops serve complementary roles - thumbnail tokens capture holistic context while local crop tokens offer fine-grained details, enabling thumbnails to guide holistic token importance evaluation from a global perspective. Through global context, visual tokens from different crops exhibit varying degrees of informativeness and visual redundancy, necessitating differentiated compression intensity to minimize information loss.
@@
Figure 2: Effects of global thumbnail and local crops. Performance (%) denotes relative scores vs. vanilla LLaVA-NeXT-7B.
@@
Building on these observations, we propose “Global Compression Commander” (i.e., GlobalCom2), which follows a “global-to-local” guided token compression design philosophy, tailored for HR-LVLMs with dynamic tiling strategy. As shown in Figure 1, GlobalCom2 first leverages holistic information from global thumbnail perspective to evaluate the information richness of each crop, adaptively allocating compression budgets across different crops. It then performs token compression for local crops through comprehensive token evaluation from both global and local perspectives, achieving differentiated compression degrees across regions. Through this “global-to-local” design, GlobalCom2 achieves training-free adaptive token compression, which selectively preserves significant visual information while removing redundancy at both global and local perspectives for HR-LVLMs with dynamic tiling.
@@
  • We empirically analyze the hierarchical nature of dynamic tiling strategy by demonstrating the distinct and complementary roles between thumbnails and crops, as well as the inherent characteristics among local crops, leading to two key observations for token compression.
  • We propose a “global-to-local” guided token compression framework GlobalCom2, which adaptively adjusts compression intensity based on crops’ information richness and preserves semantically important tokens with fine-grained details across the entire image.
  • Extensive experiments on LLaVA-NeXT and LLaVA-OneVision demonstrate superior efficiency-performance trade-offs, maintaining 90% performance with 90% token reduction, while achieving substantial memory savings and throughput improvement.
Figure 3: Effects of different local crops. Crop importance is quantified through accumulating attention scores between thumbnail patches and the [CLS] token within each crop region.
@@ 2 Related Work
2.1 High-resolution LVLMs
LVLMs generally combine a vision encoder and a large language model (LLM) decoder, bridged by a visual projector that maps visual features into the LLM’s embedding space. Representative works include BLIP-2 [22] which utilizes a Q-Former as the visual projector, and LLaVA-1.5 [28] which employs a multi-layer perceptron (MLP) projector to integrate visual and textual information.
Previous LVLMs [28, 2] resize images to a fixed resolution before processing, which often leads to shape distortion and loss of local details in high-resolution images. To address this, a series of LVLMs have been proposed to handle high-resolution images, which can be categorized into three groups: (1) Hybrid resolution methods using dual visual encoders, represented by Mini-Gemini [24] and LLaVA-HR [34]. (2) Native resolution methods adopting NaViT-style [11] encoders, represented by Qwen2-VL [42]. (3) Dynamic tiling methods, represented by LLaVA-NeXT [29] and InternVL-1.5 [9], which split images into regions and process them with a single encoder. Due to its efficiency and effectiveness in vision encoding, dynamic tiling has been widely adopted in subsequent works [30, 6, 21].
While increasing local visual tokens enhances detail recognition and hallucination suppression, it also introduces computational challenges in inference speed and memory usage. In this work, we focus on improving the efficiency of dynamic tiling-based HR-LVLMs.
2.2 Token Compression for LVLMs
Token compression, which aims to reduce the sequence length of tokens for computation efficiency, has been widely adopted for model acceleration [37, 25, 3]. For LVLMs, recent works have focused on training-free token compression paradigms, which can be broadly categorized into two streams: (1) Token compression during Vision Encoding stage [38, 45, 16]: For example, FasterVLM [52] utilizes [CLS] attention scores from the visual encoder to re-rank and retain top visual tokens. (2) Token compression during LLM Decoding stage [8, 46, 26]: Representative works include FastV [8], which prunes tokens based on self-attention scores in LLM, and SparseVLM [53], which adaptively prunes visual tokens based on their attention scores with text tokens. However, when applying these methods to HR-LVLMs with dynamic tiling, they treat all visual tokens with uniform compression strategies, overlooking the distinct roles of global thumbnail and local crop tokens in image understanding for HR-LVLMs.
In this paper, unlike existing token compression methods, we propose a “global-to-local” guided compression strategy that leverages global thumbnail as a “commander” to guide token compression in HR-LVLMs.
3 Analysis of HR-LVLM with Dynamic Tiling
@@ 3.1 Preliminary
We take LLaVA-NeXT [29], a widely-adopted HR-LVLM with dynamic tiling, as an example to conduct our analysis.
Model Architecture. Given the input image and textual instructions, LLaVA-NeXT generates instruction-following responses in an autoregressive manner through two stages:
  • Vision Encoding: LLaVA-NeXT adopts CLIP-ViT [36] to transform raw pixels into unified visual representations, followed by a two-layer MLP Projector that aligns visual embeddings with the language embedding space.
  • LLM Decoding: The decoder-only LLM takes concatenated image and prompt tokens as input and decodes them through next-token prediction to generate responses.
@@
Dynamic Tiling Input Pre-processing. LLaVA-NeXT processes arbitrary-resolution images through adaptive dynamic cropping using seven grid templates: $\{2\times 2,1\times\{2,3,4\},\{2,3,4\}\times 1\}$. For an input image of size $W\times H$, it is scaled to $(336\times a)\times(336\times b)$ based on the selected template, producing $n=a\times b$ local crops $\mathbf{X}^{L}$ of size $336\times 336$ (with padding if necessary) and a thumbnail $\mathbf{X}^{G}$. With $N$ tokens per crop, the total token length is $(1+n)\times N$.
Visual Tokens Post-processing. Following [28], LLaVA-NeXT removes padding tokens and uses special tokens to mark image boundaries, preserving aspect ratio information while improving efficiency.
3.2 Stepping into Dynamic Tiling Mechanism
@@
HR-LVLMs process high-resolution images using dynamic tiling with thumbnails and crops. We analyze their characteristics to guide token compression design.
@@
(I) Functions of Thumbnail and Crops. We begin with an exploratory experiment to investigate how thumbnail and crops contribute to image understanding in HR-LVLMs by using them separately as input to LLaVA-NeXT-7B model.
@@
As shown in Figure 2, using global thumbnail alone yields superior performance on general visual perception benchmarks such as SQA [33], MME [12], MMBench [32], and MM-Vet [48]. This advantage stems from global thumbnails providing holistic visual information through complete ViT encoding. In contrast, using only local crops, where ViT independently encodes each crop, shows inferior performance on these general visual perception tasks. However, local crops excel in fine-grained perception tasks like TextVQA [39] (VQAT, testing text-centric visual understanding) and POPE [23] (testing model hallucination) by providing detailed features. Therefore, we identify that:
@@
Observation ❶: Global thumbnails and local crops serve complementary functions in HR-LVLMs with dynamic tiling: the former acts as a “comprehensive visual extractor” for holistic representations, while the latter serves as a “detailed visual capturer” for fine-grained representations.
Figure 4: Overall framework of GlobalCom2. GlobalCom2 guides token compression for HR-LVLMs through: 1) compressing thumbnail tokens (blue path); and 2) compressing crop tokens (yellow paths) by: (a) adaptively adjusting compression intensity based on visual information richness at global view, and (b) performing compression according to token informativeness at global and local views.
@@
(II) Information Richness in Different Crops. Given that dynamic tiling leads to distinct visual representations across crops, we investigate the variance through both qualitative and quantitative perspectives.
Through visual inspection of input image in Figure 4, we observe that the upper two crops contain significant visual content (e.g., football players), while the lower two crops exhibit considerable visual redundancy (predominantly grass areas). To quantitatively validate this observed variation in informational content, we leverage insights from previous works [25, 16], which demonstrated that attention scores between [CLS] and patch tokens effectively reflect token importance. As shown in the upper middle of Figure 4, our visualization of attention scores between patch tokens and the [CLS] token from CLIP-ViT’s last layer reveals notably higher attention values in the upper regions, confirming richer semantic information in upper crops from a global perspective.
Building upon this qualitative analysis, we conduct quantitative studies to validate our observations. Since attention scores with the [CLS] token effectively measure token-level semantic richness, we leverage these scores to assess crop importance by by computing the sum of partitioned attention scores within each crop region. To examine crop contributions in HR-LVLMs, we conduct experiments on LLaVA-NeXT-7B by selectively dropping local crops (i.e., dropping the most/least important one). As shown in Figure 3, dropping the most versus least important crops leads to significant performance gaps across visual understanding tasks, with an average drop of 2.4% across six benchmarks and particularly notable in VQAT (5.2% gap). Based on the above analysis, we identify that:
@@
Observation ❷: Local crops exhibit varying information richness, leading to different contributions to the overall visual understanding of HR-LVLMs with dynamic tiling, with visually informative crops being particularly crucial for capturing fine-grained local details.
@@ 4 Global Compression Commander
As discussed in Section 3, HR-LVLMs’ dynamic tiling strategy leads to significant token length increase ( in LLaVA-NeXT [29], 10× in LLaVA-OneVision [21]), where the quadratic complexity of self-attention [41] becomes a major bottleneck. To address this challenge, we aim to directly improve HR-LVLMs’ inference speed through visual token compression, as illustrated in Figure 4.
Motivated by observation in Section 3.2 that global thumbnail serves as a “comprehensive visual extractor” providing visual representations from a holistic perspective, we leverage it as a “Global Compression Commander” (GlobalCom2) to guide token compression for HR-LVLMs, as illustrated in Figure 4. In what follows, we first elaborate on how our GlobalCom2 guides token compression for the global thumbnail.
Since visual token compression aims to preserve informative tokens while eliminating redundant representations in visual content, the key for global thumbnails lies in identifying tokens that capture essential holistic information. Given that the [CLS] token serves as an effective global image representation [25, 16], GlobalCom2 leverages the last ViT layer’s attention map to compute average attention between each global thumbnail token and the [CLS] token (blue path in Figure 4). Specifically, for 1D token sequence $\mathbf{X}^{G}$ of length $N$ in the global thumbnail, the importance score ${s}_{i}^{G}$ for the $i$-th token is:
@@
where $\mathbf{q}^{\texttt{CLS}}$ and $\mathbf{K}\in\mathbb{R}^{N\times D}$ are query of [CLS] and keys of $\mathbf{X}^{G}$. Based on these scores, tokens with higher ${s}^{G}$ capture richer semantic information and should be preserved. Given a preset token retention ratio $R$ (%), GlobalCom2 preserves the top-$k$ ($k=R\times N$) tokens ranked by ${s}^{G}$:
$\mathbf{X}^{G}\rightarrow\mathbf{\hat{X}}^{G}=\text{TopK}(\mathbf{X}^{G},s^{G},R\times N)$ (2)
where $\mathbf{\hat{X}}^{G}$ is the compressed thumbnail sequence. This training-free compression mechanism reduces computational cost by preserving only the informative tokens.
Due to dynamic tiling’s multi-partitioning process, local crops produce multiple independent token sequences containing distinct visual contents, with each crop exhibiting different levels of redundancy and collectively leading to longer sequences compared to the global thumbnail. Inspired by observation , local crops with different semantic richness warrant different token compression degrees. Specifically, semantically rich crops should retain more tokens to preserve crucial local details, while less informative ones can be more aggressively compressed to help LLMs focus on informative visual signals.
To implement such adaptive compression, motivated by observations -, we aim to guide the token compression of local crops from a global perspective. Specifically, GlobalCom2 employs two aspects of global guidance: (a) Adaptive Compression Adjustment that dynamically adapts compression intensity for each crop based on its information richness, and (b) Holistic Token Evaluation that assesses token informativeness from both local and global perspectives. We detail these two global guidance below.

### v5

@@ Abstract
Large vision-language models (LVLMs) excel at visual understanding, but face efficiency challenges due to quadratic complexity in processing long multi-modal contexts. While token compression can reduce computational costs, existing approaches are designed for single-view LVLMs and fail to consider the unique multi-view characteristics of high-resolution LVLMs with dynamic cropping. Existing methods treat all tokens uniformly, but our analysis reveals that global thumbnails can naturally guide the compression of local crops by providing holistic context for informativeness evaluation. In this paper, we first analyze dynamic cropping strategy, revealing both the complementary nature between thumbnails and crops, and the distinctive characteristics across different crops. Based on our observations, we propose “Global Compression Commander” (i.e., GlobalCom2), a novel plug-and-play token compression framework for HR-LVLMs. GlobalCom2 leverages thumbnail as the “commander” to guide the compression of local crops, adaptively preserving informative details while eliminating redundancy. Extensive experiments show that GlobalCom2 maintains over 90% performance while compressing 90% visual tokens, reducing FLOPs and peak memory to 9.1% and 60%.
Code: https://github.com/xuyang-liu16/GlobalCom2.
@@ 1 Introduction
By bridging visual encoders with large language models (LLMs) (Touvron et al. 2023; Yang et al. 2024), large vision-language models (LVLMs) (Liu et al. 2024a; Bai et al. 2023) have recently achieved remarkable progress. As LVLMs advance towards high-resolution image understanding (HR-LVLMs), dynamic cropping has emerged as a de facto standard, represents a high-resolution image as a global thumbnail combined with a set of local crops. While this enables models like LLaVA-NeXT (Liu et al. 2024b) and InternVL 3 (Zhu et al. 2025) to capture fine-grained details with vision encoding efficiency, it also introduces challenges with increased visual tokens and hierarchical visual context.
To enhance the efficiency of LVLMs, recent efforts have increasingly adopted token compression approaches (Cha et al. 2024; Shang et al. 2025), which reduce visual tokens while preserving essential information. These architecture-agnostic methods achieve optimal efficiency-accuracy trade-offs and have proven effective for LVLM acceleration. However, these methods were primarily designed for traditional single-view architectures (the entire image in Figure 1 top). While the multi-view dynamic cropping approach enables more fine-grained visual understanding, it further increases the number of visual tokens and computational overhead.
Figure 1: Design philosophy of “global-to-local” guided token compression. GlobalCom2 evaluates the information richness of local crops from a global perspective to preserve informative regions while removing redundant ones.
@@
Figure 2: Complementary roles of global thumbnail and local crops in HR-LVLMs with dynamic cropping. Performance (%) denotes relative scores of LLaVA-NeXT-7B.
However, directly applying existing token compression methods to HR-LVLMs overlooks three critical issues: (i) Global context neglect: Current methods disregard the crucial role of global thumbnails in extracting holistic context and guiding visual understanding (Figure 2). (ii) Information richness disparity: They fail to account for varying information density across different crops (Figure 3), resulting in performance gaps exceeding 5% on high-resolution tasks when comparing the removal of most versus least informative crops. (iii) Content-agnostic positional bias: Inner-LLM question-aware compression methods (Chen et al. 2024b; Xing et al. 2025; Zhang et al. 2025) systematically allocate more tokens to later-positioned crops irrespective of their actual informative content (Figure 4), inducing severe multi-modal hallucinations under extreme compression (Table 1), with detailed analysis provided in Appendix (Figure 9). These three oversights collectively result in over-compression of information-rich regions and disruption of the visual-semantic hierarchy (Figure 7).
To bridge this critical gap, we first conduct systematic analysis of dynamic cropping in Section 3, identifying two key observations: Thumbnail and crop tokens serve complementary roles - thumbnails capture holistic context while crops provide fine-grained details, enabling global importance evaluation. Through global context, tokens from different crops exhibit varying informativeness, requiring differentiated compression to minimize information loss.
@@
Building on these observations, we propose “Global Compression Commander” (GlobalCom2), which follows a “global-to-local” guided token compression philosophy tailored for dynamic cropping-based HR-LVLMs. In Figure 1, GlobalCom2 leverages holistic information from thumbnails to evaluate each crop’s information richness, adaptively adjusting compression intensity. It performs token compression through comprehensive evaluation from both global and local perspectives, achieving differentiated compression across regions while preserving significant information. This training-free approach extends seamlessly to video large language models (VideoLLMs) for efficient video understanding and can be integrated with existing question-aware methods. Notably, integrating this “global-to-local” design with FastV (Chen et al. 2024b) and SparseVLM (Zhang et al. 2025) achieves significant improvements of 5.3% and 5.2% across benchmarks while alleviating multi-modal hallucination caused by positional bias.
@@
  • Systematic Dynamic Cropping Analysis: We empirically analyze the hierarchical nature of dynamic cropping and thoroughly investigate existing token compression methods for HR-LVLMs, identifying fundamental causes of critical information over-compression.
  • Global-to-Local Compression Philosophy: We propose GlobalCom2, a training-free framework that adaptively adjusts compression intensity based on crop information richness assessment and preserves semantically important tokens across the entire visual hierarchy.
  • Superior Performance-Efficiency Trade-offs: Extensive experiments on multiple HR-LVLMs and VideoLLMs demonstrate exceptional balance of GlobalCom2, maintaining over 90% performance with 90% token reduction, while achieving substantial memory savings and throughput improvements.
Figure 3: Varying contributions of local crops. Importance is quantified by the accumulated attention scores between thumbnail patches and [CLS] token within each crop.
@@ 2 Related Work
High-resolution LVLMs. Large vision-language models (LVLMs) integrate vision encoders and LLM decoders via projectors for feature alignment. Early LVLMs (Liu et al. 2024a; Bai et al. 2023) resize images to fixed resolutions, causing shape distortion and detail loss. To address this, high-resolution LVLMs have emerged in three categories: (i) Hybrid resolution methods using dual visual encoders (Li et al. 2024c; Luo et al. 2025). (ii) Native resolution methods with NaViT-style encoders (Wang et al. 2024; Guo et al. 2025). (iii) Dynamic cropping methods (Li et al. 2025; Zhu et al. 2025), which split images into regions for single-encoder processing. Dynamic cropping has gained widespread adoption (Lu et al. 2025; Chen et al. 2025) due to its vision encoding efficiency. However, increased visual tokens introduce computational challenges in inference speed and memory usage. Our work focuses on improving the efficiency of dynamic cropping-based HR-LVLMs.
Token Compression for LVLMs. Token compression has been widely adopted for LVLM acceleration through training-free approaches at two stages: (i) Vision Encoding stage (Shang et al. 2025; Yang et al. 2025): FasterVLM (Zhang et al. 2024) leverages [CLS] attention scores to prune visual tokens. (ii) LLM Pre-filling stage (Xing et al. 2025; Ye et al. 2025): FastV (Chen et al. 2024b) prunes tokens based on LLM self-attention, while SparseVLM (Zhang et al. 2025) uses cross-modal attention scores. Some methods combine both stages (Liu et al. 2024c). However, these methods treat all visual tokens equally in a “flat” token space, ignoring HR-LVLMs’ hierarchical structure where thumbnails and crops serve distinct roles.
Our work is the first to systematically quantify this oversight’s severe consequences, revealing critical failure modes including uniform over-compression and positional bias. We address this by proposing a “global-to-local” compression framework that leverages hierarchical structure, representing a shift toward structure-aware token compression.
3 Analysis of Dynamic Cropping
@@ 3.1 Preliminary
Figure 4: Content-agnostic positional bias. LLM attention-guided methods (e.g., FastV) assign higher scores (bars) to later tokens, regardless of their content or input order (second row: sequential crops; third row: reversed crops).
@@
We take LLaVA-NeXT, a widely-adopted HR-LVLM with dynamic cropping, as an example to conduct our analysis.
Model Architecture. Given input image and text, LLaVA-NeXT generates responses through: (i) Vision Encoding: ViT (Radford et al. 2021) converts pixels to embeddings via MLP projector. (ii) LLM Decoding: LLM processes concatenated tokens, generating responses auto-regressively.
@@
Pre-processing. LLaVA-NeXT uses grid templates: $\{2\times 2,1\times\{2,3,4\},\{2,3,4\}\times 1\}$. Images of size $W\times H$ are scaled to $(336\times a)\times(336\times b)$, yielding $n=a\times b$ local crops $\mathbf{X}^{L}$ and a thumbnail $\mathbf{X}^{G}$. Total token length is $(1+n)\times N$.
Post-processing. LLaVA-NeXT removes padding and adds boundary tokens to preserve aspect ratios.
3.2 Stepping into Dynamic Cropping
@@
HR-LVLMs process high-resolution images using dynamic cropping with thumbnails and crops. We analyze their characteristics to guide token compression design.
@@
(I) Functions of Thumbnail and Crops. We investigate how thumbnail and crops contribute to image understanding in HR-LVLMs through experiments with LLaVA-NeXT-7B.
@@
Figure 2 shows that thumbnail alone performs better on general visual perception tasks (e.g., SQA (Lu et al. 2022) and MMBench (Liu et al. 2024d)), benefiting from holistic visual information. In contrast, local crops excel in fine-grained tasks like text understanding (TextVQA (Singh et al. 2019)) and hallucination detection (POPE (Li et al. 2023)) by providing detailed features. Therefore, we identify that:
@@
Observation ❶: Global thumbnails and local crops serve complementary functions in HR-LVLMs with dynamic cropping: the former acts as a “comprehensive visual extractor” for holistic representations, while the latter serves as a “detailed visual capturer” for fine-grained representations.
@@
(II) Information Richness in Different Crops. We investigate visual representation variance across crops through qualitative and quantitative analysis.
Figure 5 shows upper crops contain rich content (e.g., football players) while lower crops show mainly grass. CLIP-ViT’s last-layer attention distribution confirms higher semantic information in upper regions.
Quantitative experiments measuring crop importance via [CLS] attention scores and drop-crop tests on LLaVA-NeXT-7B (Figure 3) reveal substantial performance gaps: averaging 2.4% across benchmarks and reaching 5.5% and 5.2% on high-resolution tasks MM-Vet (Yu et al. 2024) and TextVQA (Singh et al. 2019) when dropping most versus least important crops. We thus identify:
@@
Observation ❷: Local crops exhibit varying information richness in global context, leading to different contributions to the overall visual understanding of HR-LVLMs with dynamic cropping, with visually informative crops being particularly crucial for capturing fine-grained local details.
Figure 5: Overall framework. GlobalCom2 guides token compression for HR-LVLMs through: 1) compressing thumbnail tokens (blue path), and 2) compressing crop tokens (yellow paths) by (a) adaptively adjusting compression intensity based on global visual richness, and (b) performing compression according to token informativeness from global and local perspectives.
@@ 4 Global Compression Commander
HR-LVLMs with dynamic cropping increase token length, making self-attention’s quadratic complexity a bottleneck. Inspired by human vision’s process of grasping scene gist before focusing on details, we propose “Global Compression Commander” (GlobalCom2) which implements a “global-to-local” compression strategy.
In Figure 5, for thumbnail token compression, we identify tokens with essential holistic information. Given the [CLS] token’s effectiveness as global image representation (Liang et al. 2022), GlobalCom2 uses the last ViT layer’s attention map to compute attention between each token and [CLS] [^1] For HR-LVLMs without [CLS] token, we propose an alternative token informativeness evaluation metric. See Appendix. (blue path in Figure 5). For the 1D token sequence $\mathbf{X}^{G}$ of length $N$, the importance score ${s}_{i}^{G}$ for the $i$-th token is:
@@
where $\mathbf{q}^{\texttt{CLS}}$ and $\mathbf{K}\in\mathbb{R}^{N\times D}$ are query of [CLS] and keys of $\mathbf{X}^{G}$. Given a preset retention ratio $R$ (%), GlobalCom2 preserves the top-$k$ ($k=R\times N$) tokens ranked by ${s}^{G}$:
$\mathbf{X}^{G}\rightarrow\mathbf{\hat{X}}^{G}=\text{TopK}(\mathbf{X}^{G},s^{G},R\times N).$ (2)
While thumbnail compression focuses on preserving holistic context, crop compression faces more complex challenges due to varying information densities across different crops. Following observation , semantically rich crops should preserve more tokens for crucial details, while less informative ones can be compressed more aggressively.
GlobalCom2 leverages the comprehensive visual knowledge from thumbnails to guide crop compression through a decoupled two-stage process: (a) Adaptive Compression Adjustment, which dynamically adapts the compression intensity for each crop based on its information richness, and (b) Holistic Token Evaluation, which assesses the informativeness of tokens from both local and global views.